Complete the ordered pair to make a solution to the given equation: (-7/2, ?); the equation is -2x+3y=25
Answer by Prithwis(166)

(-7/2, ?); the equation is -2x+3y=25
x=-7/2
-2(-7/2)+3y=25
3y=25-7
y=18/3
y=6
(-7/2,6)